OpenStudy (anonymous):

some one plzz help Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (anonymous):

multiply the first equation by 2, then subtract the second equation from it.

OpenStudy (igreen):

Solve for \(x\) in the first equation: \(x + 3y = -1\) \(x = -3y - 1\)

OpenStudy (anonymous):

ok so the second equation would becom 2(-3y-1) + 2y=6

OpenStudy (igreen):

Now we can plug in \(-3y-1\) for \(x\) in the second equation: \(2(-3y-1) + 2y = 6\) Solve for \(y\): \(-6y - 2 + 2y = 6\) \(-6y + 2y = 8\) \(-4y = 8\) \(y = -2\)

OpenStudy (igreen):

Yes, that's right. @elketen

OpenStudy (anonymous):

thanks a lot

OpenStudy (igreen):

Now we have our y-value, -2. Then we can plug it into any of the equations: \(x + 3y = -1\) \(x + 3(-2) = -1\) \(x - 6 = -1\) \(x = 5\)

OpenStudy (igreen):

So the solution is \((5,-2)\).

OpenStudy (anonymous):

ok i get it thanks i was confuse

OpenStudy (igreen):

No problem :)

OpenStudy (anonymous):

can you help me with another one

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

No problem. Just close this and open another one.